in a circle arc ab=3, arcbc=5, arcca=10. then angle a,b,c are
 

Grade:11

1 Answers

DR STRANGE
7 years ago
HI ISHITA,
WELL THE QUESTION IS NOT CLEAR ,WHAT ARE THE \angle A ,\angle B ,\angle C?
IF YOU MEAN THE ANGLES MADE AT THE CENTER OF THE CIRCLE THEN HERE IS THE SOLUTION,
JUT SUM THE LENTHS OF ARCS, 3+5+10=18,
NOW FIND RATIO FOR EACH ARC,
FOR ARC AB ,=3/18
FOR ARC BC, =5/18
FOR ARC CA, =10/18
WE KNOW SUM OF ALL ANGELS MADE AT CENTER OF THE CIRCLE IS 3600.
SO \angle A= (3/18)\times360 \angle A= (3/18)\times360 = 60
HENCE YOU CAN FIND \angle B AND \angle C BY JUST MULTIPLYING THE RATIO WITH 3600.
HAVE A GOOD DAY :)